A 50-gram block rests on an inclined plane 15o with the horizontal. The coefficient of friction is 0.10. Find:  a) the acceleration of the block, b) the work done against friction 2 minutes after starting from rest, and c) its average power.

 


1
Expert's answer
2021-11-29T15:13:57-0500

a) Let's apply the Newton's Second Law of Motion:


"mgsin\theta-\mu mgcos\theta=ma,""a=g(sin\theta-\mu cos\theta),""a=9.8\ \dfrac{m}{s^2}\times(sin15^{\circ}-0.1\times cos15^{\circ})=1.59\ \dfrac{m}{s^2}."

b) Let's first find the final velocity of the block 2 minutes after starting from rest:


"v=v_0+at=0+1.59\ \dfrac{m}{s^2}\times120\ s=190.8\ \dfrac{m}{s}."

Then, we can find the distance traveled by the block during this time:


"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\times(0+190.8\ \dfrac{m}{s})\times120\ s=11448\ m."

Finally, we can find the work done against friction 2 minutes after starting from rest:


"W_{fr}=F_{fr}d=\mu mgdcos\theta,""W_{fr}=0.1\times0.05\ kg\times9.8\ \dfrac{m}{s^2}\times11448\ m\times cos15^{\circ}=542\ J."

c) By the definition of the power, we get:


"P=\dfrac{W}{t}=\dfrac{542\ J}{120\ s}=4.52\ W."